OpenStudy (anonymous):

Find unit vector a-same direction as v b-opposite direction of v v=(2,2)

OpenStudy (anonymous):

could you help me find the unit vector?

OpenStudy (anonymous):

divide each component by the length of v to get the unit vector.

OpenStudy (anonymous):

hmm...?

OpenStudy (anonymous):

how would I do that?

OpenStudy (anonymous):

what's the length or norm of v?

OpenStudy (anonymous):

well, the coordinates are at (2,2)

OpenStudy (anonymous):

I know that much

OpenStudy (anonymous):

\[||v|| =\sqrt{2^2+2^2}=2\sqrt{2}\] unit vector has length of 1 => \[\frac{v}{||v||}\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

how did you find out that the unit vector has a length of 1 based on just achieving the result 2 and the square root of 2?

OpenStudy (anonymous):

a unit vector is defined to have a lenght of 1... similar to the unit circle: defined as having a radius of 1.

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

gotta know your defs for math! it's a language

OpenStudy (anonymous):

what's the next step?

OpenStudy (anonymous):

I got llvll=2 and the square root of 2 and the unit vector has a length of 1

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

sorry if I am clueless, I just have a different precalculus book than my professor

OpenStudy (anonymous):

once I grasp the logic of the problem, it is like cracking a lock open

OpenStudy (anonymous):

to find the unit vector which points in the direction of v, simply divide v by its norm. this involves dividing each compnent of v by its norm. \[\vec{u}=\frac{\vec{v}}{||\vec{v}||}=\left(\frac{2}{2\sqrt{2}},\,\frac{2}{2\sqrt{2}}\right) \]

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

to find the one pointing in the opposite direction, simply multiply the unit vector by -1. Also, you should simplify the previous.

OpenStudy (anonymous):

so the opposite direction would be \[(-2/2\sqrt{2}, -2/2\sqrt{2}\]

OpenStudy (anonymous):

yes, but simplify. no radicals on the bottom.

OpenStudy (anonymous):

(-2/2 square root of 2, -2/2 square root of 2

OpenStudy (anonymous):

-2 square root of 2/4?

OpenStudy (anonymous):

learn to use the equation editor... it makes things much clearer! \[\vec{u}=\left(\frac{\sqrt{2}}{2},\,\frac{\sqrt{2}}{2}\right)\] \[-\vec{u}=\left(-\frac{\sqrt{2}}{2},\,-\frac{\sqrt{2}}{2}\right)\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

you're welcome